OpenStudy (anonymous):

write the improper fraction 31/20 as a mixed number?

OpenStudy (vijay):

1 (11/ 20)

OpenStudy (anonymous):

is the fraction 18 11/20 in simplest form?

jimthompson5910 (jim_thompson5910):

\[\frac{31}{20}=\frac{20+11}{20}=\frac{20}{20}+\frac{11}{20}=1+\frac{11}{20}=1\frac{11}{20}\] So \[\frac{31}{20}=1\frac{11}{20}\]